Hello tutors , can you answer this problem for me . If I subtract from twice my present age three times my age six years ago , i get my present . What is my age ? tanx so much
Answer by ankor@dixie-net.com(22740)

If I subtract from twice my present age, three times my age six years ago,
I get my present . What is my age ?
Let a = your present age
Just write an equation for what it says:
a = 2a - 3(a-6)
a = 2a - 3a + 18
a = -a + 18
a + a = 18
2a = 18
a = 9 yrs is your present age
:
:
:
Check that in the given statement:
2(9) - 3(9-6) =
18 - 3(3) = 9